OpenStudy (anonymous):

What is 4/9 / 1/3 A. 1 1/3 B. 1 2/3 C. 4/24 D. 7/9

OpenStudy (k_lynn):

what is that @AidanWithanAnota

OpenStudy (ilovewolf):

4/9 / 1/3 = ? = ( 4 * 3) / ( 1 * 9) = 12/9 = 4/3 =1 1/3 so the answer would be A. c: